Blue Shade wants to launch a newproduct called Shady Blue in the market.The sales manager needs to presentthe opportunity to management. Heapproaches you to assist him in calculating the required information.

He providesyou with thefollowinginformation.

 

Purchase price of theproduct

R125,50/unit

 

Packaging cost

R10,00/unit

 

Labour needed towrap the productbefore it can

be delivered. (Theproduct must bewrapped and cannot be soldwithout thewrapping)

Wrap 2products

per hour.The employeeswill be paid R160 per day. The companywork a 8hour day.Theemployeeswill beutilised somewhereelse f thereare no products tobe wrapped

 

A supervisor needs to beappointed at amonthly cost of R15,000.

 

 

Delivery cost to thewholesalers will be charged at

R300 per 10 unitsdelivered.

 

 

Additional space will be rented at R5,000per

month.

 

 

Additional general administrationexpenses will

amount to R2,500per month

 

 

PBA4807

 

Required:

 

Assist the sales manager in calculatingthe following:

 

1.    The estimated sales price per unit. The company’s policy is a mark-up of 65% on variable cost.                                                                                                                               (5)

2.    The contribution per unit.                                                                                               (1)

 

3.    Break-even units to be sold to cover the additional costs.                                            (1)

 

4.    The number of units to be sold to achieve a profit before tax of 20% ofthe sales value.

(2)

5.    The number of units to be sold to achieve a profit after tax of 15% of the sales value. The tax rate is 28%.                                                                                                              (3)
Answer #1

Variable Costs:

Raw material cost = R125.50 per unit

Packaging cost per unit = R 10 per unit

Direct labour cost = R 160 per day. It takes 1 hr to wrap 2 units, hence in 8 hrs per day, 2* 8 = 16 products could be wrapped.So wrapping 16 products wrapping costs R 160, so wrapping cost for one unit = 160/16 = R10 per unit.

Delivery cost = 30 per unit

Total Variable Cost per unit = 125.50+10+10+30 = R175.50

Total Fixed Cost = Rentnal cost+ Admin Cost+Manager's Salary = 5000+2500+15000 = R22,500

Total Cost = Fixed cost + Variable Cost

Now the number of units produced in a month is not given. so let us assume there is a 100 unit production in a month.

Variable cost = 175.5 * 100 = 17,550

Fixed cost = 22500

Hence total cost = 17550+22500 = R40,050

1. Sales price per unit with a 65% markup

= Total cost + 65% of variable cost = 40050+ 65% of 17550 = R51457.50

Sales price per unit = R51457.50/100 = R514.575

2. Contribution = Unit sales price - Unit variable cost = R514.575 - 175.50 = R339.075

3.Break even point = Fixed Cost/(sales pere unit- Variable cost) = 22500/(514.575-175.5) = 66.357 or approx 66 units.

4. Profit = sales - total cost = (sales price per unit * number of units sold) - (variable cost * no. of units sold) - fixed cost

lets units sold = s

20/100*514.575*s = 514.575s-175,50s-22500

=> k = 22500/235.50 = 96 units

Therefore in order to earn desired profit sales needed is 96 units.